# Masked Autoencoders are Parameter-Efficient Federated Continual Learners

### v2

@@ Abstract
Federated learning is a specific distributed learning paradigm in which a central server aggregates updates from multiple clients’ local models, thereby enabling the server to learn without requiring clients to upload their private data, maintaining data privacy. While existing federated learning methods are primarily designed for static data, real-world applications often require clients to learn new categories over time. This challenge necessitates the integration of continual learning techniques, leading to federated continual learning (FCL). To address both catastrophic forgetting and non-IID issues, we propose to use masked autoencoders (MAEs) as parameter-efficient federated continual learners, called pMAE. pMAE learns reconstructive prompt on the client side through image reconstruction using MAE. On the server side, it reconstructs the uploaded restore information to capture the data distribution across previous tasks and different clients, using these reconstructed images to finetune discriminative prompt and classifier parameters tailored for classification, thereby alleviating catastrophic forgetting and non-IID issues on a global scale. Experimental results demonstrate that pMAE achieves performance comparable to existing prompt-based methods and can enhance their effectiveness, particularly when using self-supervised pre-trained transformers as the backbone. Code is available at: https://github.com/ycheoo/pMAE.
@@ I Introduction
Building on the insights from Zhai et al. [45], we propose to use masked autoencoders (MAEs) as parameter-efficient federated continual learners, called pMAE. Originally introduced as scalable vision learners, MAEs [13] leverage self-supervised learning [4, 13, 53] to learn effective visual representations. Instead of focusing solely on representation learning, pMAE emphasizes the ability of MAEs to reconstruct images from restore information. In pMAE, client-side learning not only prioritizes classification as the primary task but also focuses on reconstruction as a secondary task. On the server side, the images reconstructed from uploaded restore information are utilized to finetune model parameters tailored for classification. Reconstructed images represent distributions from old tasks and diverse client data, thereby enabling the global mitigation of catastrophic forgetting and non-IID issues.
In FCL, the limitations of client-side computational resources make it burdensome to train a transformer-based model (such as MAEs) [34, 7, 9] from scratch; therefore, we employ pre-trained transformers [12] as the backbone. To further enhance parameter efficiency, we freeze the parameters of the pre-trained transformer and leverage prompt tuning [21, 17]. Specifically, clients utilize discriminative prompt and classifier parameters for classification while employing reconstructive prompt for reconstructing masked images. After completing training, clients randomly extract restore information from a subset of the images and upload them to the server. The server then reconstructs images using reconstructive prompt and restore information, finetuning discriminative prompt and classifier parameters in the process, which mitigates performance degradation from catastrophic forgetting and non-IID issues on a global scale. Given the generality of our approach, it can be directly integrated with existing prompt-based methods [39, 38, 31] to enhance their performance.
@@
- We introduce prompt tuning as a parameter-efficient method for finetuning pre-trained transformers, employing discriminative prompt for the classification task and reconstructive prompt for the reconstruction task.
@@ III-B Masked Autoencoders
Figure 2: The overview of proposed pMAE framework. 1) Client: Extract labeled restore information using Eq. (3) and optimize prompts and classifier parameters of the local model by $\mathcal{L}_{client}$. 2) Server: Generate reconstructed images using Eq. (4) and finetune discriminative prompt and classifier parameters of the aggregated global model by $\mathcal{L}_{server}$. 3) Lightweight prompts, classifier parameters and labeled restore information are transmitted between clients and the server, while no additional data is stored on each client.
The MAE encoder is a Vision Transformer (ViT) model [9], where input images are divided into non-overlapping patches, and a significant portion (e.g., 75 %) of these patches are randomly masked. The remaining unmasked patches are encoded into visible tokens by the encoder and passed to the decoder along with restore ids, which indicate the positions of masked tokens.
@@ IV Methodology of the pMAE
On the server side, the received restore information and reconstructive prompt are used to reconstruct images, which are subsequently used to finetune the discriminative prompt and classifier parameters, mitigating catastrophic forgetting and non-IID issues from a global perspective.
@@ IV-A Client Side: Learning from Prompting
IV-B Server Side: Finetuning from Reconstructing
@@
This reconstructed global dataset is then used to finetune the discriminative prompt and classifier parameters as:
@@
During the subsequent server finetuning, if the current task is still ongoing, ($\left(\{{\mathcal{R}}_{k}\}_{k=1}^{K},\bm{p}_{r}\right)$ is utilized for image reconstruction. If the current task has ended, the aggregated labeled restore information and reconstructive prompt are merged into the restore pool $\mathcal{P}_{g}$. We then use $\mathcal{P}_{g}=\left\{\left(\{{\mathcal{R}}_{k}^{i}\}_{k=1}^{K},\bm{p}_{r}^{i}\right)\right\}_{i=1}^{t}$, which contains both old task data and current task data, to generate a reconstructed global dataset $\hat{\mathcal{D}}_{g}$ and use Eq. (6) to finetune the discriminative prompt and classifier parameters.
@@ V-A Experimental Setup
Implementation Details. The Adam optimizer [19] is applied across all methods, utilizing a consistent batch size of 128 and a learning rate of $lr=1e^{-3}$. The number of clients is set to $K=10$, with a total of $R_{all}=200$ communication rounds and local update epochs $E=5$. For each task, communication rounds are determined as $R={\textstyle{\frac{R_{all}}{T}}}$. For pMAE, the uploaded number of restore information is set to $u=4$, and finetuning on the server is performed over $E_{server}=5$ epochs.
@@ V-B Performance Comparison
Although the out-of-distribution data in ImageNet-R limits the effectiveness of the discriminative prompt, the mechanism of using restore information for reconstruction on the server to finetune the discriminative prompt and classifier parameters enables pMAE to effectively mitigate performance degradation caused by catastrophic forgetting and non-IID issues, allowing it to achieve optimal performance in other settings. Furthermore, as the non-IID degree increases, the performance gap between pMAE and other methods becomes more significant, consistent with the observations from CUB-200.
@@ V-C Enhancing Other Methods with pMAE
The generality of pMAE enables seamless integration with existing prompt-based methods [39, 38, 31] to enhance performance. Since these methods already utilize an encoder for feature extraction, integrating pMAE simply requires adding a decoder to establish a complete MAE. During client training, the reconstructive prompt is optimized with MSE loss (as outlined in Section IV). Restore information is then extracted on the client, uploaded to the server, and used with the reconstructive prompt for reconstruction. These reconstructed images are subsequently used for finetuning parameters tailored for classification, without any further modifications.
@@ V-D Ablation Study
Specifically, the ablation of reconstructive prompt results in a decline in the quality of image reconstructed on the server side. As shown in Figures 9 and 10, image reconstructed with reconstructive prompt exhibit superior quality, which is crucial for the finetuning process on the server side. The reconstructive prompt offers a more significant quality improvement for the Sup-based model compared to the iBOT-based model, which can also be confirmed by the performance degradation resulting from ablating reconstructive prompt in Tables III and IV.

### v3

@@ Abstract
Federated learning is a specific distributed learning paradigm in which a central server aggregates updates from multiple clients’ local models, thereby enabling the server to learn without requiring clients to upload their private data, maintaining data privacy. While existing federated learning methods are primarily designed for static data, real-world applications often require clients to learn new categories over time. This challenge necessitates the integration of continual learning techniques, leading to federated continual learning (FCL). To address both catastrophic forgetting and non-IID issues, we propose to use masked autoencoders (MAEs) as parameter-efficient federated continual learners, called pMAE. pMAE learns reconstructive prompt on the client side through image reconstruction using MAE. On the server side, it reconstructs the uploaded restore information to capture the data distribution across previous tasks and different clients, using these reconstructed images to fine-tune discriminative prompt and classifier parameters tailored for classification, thereby alleviating catastrophic forgetting and non-IID issues on a global scale. Experimental results demonstrate that pMAE achieves performance comparable to existing prompt-based methods and can enhance their effectiveness, particularly when using self-supervised pre-trained transformers as the backbone. Code is available at: https://github.com/ycheoo/pMAE.
@@ I Introduction
Building on the insights from Zhai et al. [45], we propose to use masked autoencoders (MAEs) as parameter-efficient federated continual learners, called pMAE. Originally introduced as scalable vision learners, MAEs [13] leverage self-supervised learning [4, 13, 53] to learn effective visual representations. Instead of focusing solely on representation learning, pMAE emphasizes the ability of MAEs to reconstruct images from restore information. In pMAE, client-side learning not only prioritizes classification as the primary task but also focuses on reconstruction as a secondary task. On the server side, the images reconstructed from uploaded restore information are utilized to fine-tune model parameters tailored for classification. Reconstructed images represent distributions from old tasks and diverse client data, thereby enabling the global mitigation of catastrophic forgetting and non-IID issues.
In FCL, the limitations of client-side computational resources make it burdensome to train a transformer-based model (such as MAEs) [34, 7, 9] from scratch; therefore, we employ pre-trained transformers [12] as the backbone. To further enhance parameter efficiency, we freeze the parameters of the pre-trained transformer and leverage prompt tuning [21, 17]. Specifically, clients utilize discriminative prompt and classifier parameters for classification while employing reconstructive prompt for reconstructing masked images. After completing training, clients randomly extract restore information from a subset of the images and upload them to the server. The server then reconstructs images using reconstructive prompt and restore information, fine-tuning discriminative prompt and classifier parameters in the process, which mitigates performance degradation from catastrophic forgetting and non-IID issues on a global scale. Given the generality of our approach, it can be directly integrated with existing prompt-based methods [39, 38, 31] to enhance their performance.
@@
- We introduce prompt tuning as a parameter-efficient method for fine-tuning pre-trained transformers, employing discriminative prompt for the classification task and reconstructive prompt for the reconstruction task.
@@ III-B Masked Autoencoders
Figure 2: The overview of proposed pMAE framework. 1) Client: Extract labeled restore information using Eq. (3) and optimize prompts and classifier parameters of the local model by $\mathcal{L}_{client}$. 2) Server: Generate reconstructed images using Eq. (4) and fine-tune discriminative prompt and classifier parameters of the aggregated global model by $\mathcal{L}_{server}$. 3) Lightweight prompts, classifier parameters and labeled restore information are transmitted between clients and the server, while no additional data is stored on each client. 4) Images are reconstructed into tensors and subsequently used for fine-tuning, ensuring that no real images are stored on the server.
The MAE encoder is a Vision Transformer (ViT) model [9], where input images are divided into non-overlapping patches, and a significant portion (e.g., 75 %) of these patches are randomly masked. The remaining unmasked patches are encoded into visible tokens and passed to the decoder along with restore ids, which indicate the positions of masked tokens.
@@ IV Methodology of the pMAE
On the server side, the received restore information and reconstructive prompt are used to reconstruct images, which are subsequently used to fine-tune the discriminative prompt and classifier parameters, mitigating catastrophic forgetting and non-IID issues from a global perspective.
@@ IV-A Client Side: Learning from Prompting
IV-B Server Side: Fine-tuning from Reconstructing
@@
This reconstructed global dataset is then used to fine-tune the discriminative prompt and classifier parameters as:
@@
During the subsequent server fine-tuning, if the current task is still ongoing, ($\left(\{{\mathcal{R}}_{k}\}_{k=1}^{K},\bm{p}_{r}\right)$ is utilized for image reconstruction. If the current task has ended, the aggregated labeled restore information and reconstructive prompt are merged into the restore pool $\mathcal{P}_{g}$. We then use $\mathcal{P}_{g}=\left\{\left(\{{\mathcal{R}}_{k}^{i}\}_{k=1}^{K},\bm{p}_{r}^{i}\right)\right\}_{i=1}^{t}$, which contains both old task data and current task data, to generate a reconstructed global dataset $\hat{\mathcal{D}}_{g}$ and use Eq. (6) to fine-tune the discriminative prompt and classifier parameters.
@@ V-A Experimental Setup
Implementation Details. The Adam optimizer [19] is applied across all methods, utilizing a consistent batch size of 128 and a learning rate of $lr=1e^{-3}$. The number of clients is set to $K=10$, with a total of $R_{all}=200$ communication rounds and local update epochs $E=5$. For each task, communication rounds are determined as $R={\textstyle{\frac{R_{all}}{T}}}$. For pMAE, the uploaded number of restore information is set to $u=4$, and fine-tuning on the server is performed over $E_{server}=5$ epochs.
@@ V-B Performance Comparison
Although the out-of-distribution data in ImageNet-R limits the effectiveness of the discriminative prompt, the mechanism of using restore information for reconstruction on the server to fine-tune the discriminative prompt and classifier parameters enables pMAE to effectively mitigate performance degradation caused by catastrophic forgetting and non-IID issues, allowing it to achieve optimal performance in other settings. Furthermore, as the non-IID degree increases, the performance gap between pMAE and other methods becomes more significant, consistent with the observations from CUB-200.
@@ V-C Enhancing Other Methods with pMAE
The generality of pMAE enables seamless integration with existing prompt-based methods [39, 38, 31] to enhance performance. Since these methods already utilize an encoder for feature extraction, integrating pMAE simply requires adding a decoder to establish a complete MAE. During client training, the reconstructive prompt is optimized with MSE loss (as outlined in Section IV). Restore information is then extracted on the client, uploaded to the server, and used with the reconstructive prompt for reconstruction. These reconstructed images are subsequently used for fine-tuning parameters tailored for classification, without any further modifications.
@@ V-D Ablation Study
Specifically, the ablation of reconstructive prompt results in a decline in the quality of image reconstructed on the server side. As shown in Figures 9 and 10, image reconstructed with reconstructive prompt exhibit superior quality, which is crucial for the fine-tuning process on the server side. The reconstructive prompt offers a more significant quality improvement for the Sup-based model compared to the iBOT-based model, which can also be confirmed by the performance degradation resulting from ablating reconstructive prompt in Tables III and IV.